RC Meeting, Friday, November 5, 2021
RC Meeting Information
RC Daily Meetings are at 10 am, remote connection link: https://bluejeans.com/861439920 There is a 10:00 RC meeting today, 11/03/2021!
- Energy change started later this morning than planned around 8:30
Hall C Status
- Will be in controlled access.
- Heidi will replace a UPS unit on SHMS cryo.
- Vladimir + techs will change SHMS aerogel material.
- Amy will help change SHMS angle to 8 deg prior to Hall work.
- Accelerator changing the energy.
Number of good pions is 10274
Current Run Plan
- Accelerator will change energy and push 5-pass beam into Hall for BLM and RF trip tuning over the weekend.
- After that, we can go to 3 pass and begin that physics
- Replace flow switches for (HMS?) magnet power supplies: next opportunistic access (Joe Beaufait) ( PENDING )
- HMS Q3 insulating vacuum spikes generating annoyance alarms that take minutes to recover. https://logbooks.jlab.org/entry/3930175 Noise levels on transducers look OK, but Junaid's plot indicates Q3 is burping: https://logbooks.jlab.org/entry/3932992 .
- Serious HB LHe low liquid level alarms. https://logbooks.jlab.org/entry/3933051 Always seems to start with a subtle, simultaneous increase in level in Q1/2/3. https://logbooks.jlab.org/entry/3929152 .
- Target ladder camera was identified to be bad: Joe Beaufait will order camera ( PENDING )
- Still running with no UPS on the SHMS cryo controls - Steve L. has ordered batteries and UPS. ( PENDING )
- HCCAM02 not functioning (lift required to replace-Brad is tracking this replacement).
- How the harp scans and carbon hole target should look (see links):
- carbon hole target beam centering procedure:
- Resetting the "ioc heartbeat epics communications" (start by clicking on the heart at upper left of the magnet controls screen
- Fall 2021 Shift Schedule
- Hall-C Live Status
- Hall-C Logbook
- MCC White Board